# Differentially Private Synthetic Data via Foundation Model APIs 2: Text

### v1

@@ 1 Introduction
Figure 1: Instead of finetuning LLMs with DP-SGD to generate synthetic text, our Aug-PE only requires inference APIs of LLMs. Aug-PE works with the latest open-source LLMs and API-based LLMs to generate DP synthetic text with improved utility on OpenReview dataset, where DP-SGD finetuning is either hard to implement or infeasible.
The state-of-the-art DP synthetic text approach is to finetune pretrained generative language models (LMs) on private data with DP-SGD (a DP variant of SGD (Abadi et al., 2016)(Yue et al., 2023; Kurakin et al., 2023; Mattern et al., 2022) (short-handed as DP finetune generator; see Fig. 1). Unlike non-DP ML applications, which have been greatly advanced by powerful LLMs such as GPT-4 (OpenAI, 2023b) and LLaMA (Touvron et al., 2023a, b) in a short time after they are released, the state-of-the-art DP synthetic text approaches are unfortunately still based on GPT-2.111The 175 billion-parameter GPT-3 has also been used for DP synthetic text (He et al., 2022). However, the solution is not publicly accessible as GPT-3 is proprietary. The reasons are: (1) Many powerful LLMs such as GPT-4, Claude, and Bard are only accessible through APIs. DP finetuning them is not feasible. [^2] Although standard finetuning APIs are provided for some of the models (OpenAI, 2023a), DP finetuning requires a special implementation and no model provides this custom API to date. (2) Even though some LLMs (e.g., LLaMA) are open-source, finetuning them with DP is resource-intensive and non-trivial to implement due to the need to calculate per-sample gradients (see § 2).
@@ 2 Background
DP synthetic text. To guarantee DP for private training data, one method involves using DP-SGD (Abadi et al., 2016) during model training for specific NLP tasks (Yu et al., 2022; Li et al., 2021). Alternatively, one can finetune pretrained generative language models, such as GPT-2, with private data using DP-SGD and then generate synthetic text datasets (Putta et al., 2022; Bommasani et al., 2019) (Fig. 1). Such DP synthetic texts can be employed in an arbitrary number of non-privately trained downstream tasks without increasing privacy loss. Studies by Yue et al. (2023); Mattern et al. (2022); Kurakin et al. (2023) indicate that training downstream models on DP synthetic text yields performance akin to directly training them on real data with DP, highlighting the good quality of synthetic data.
@@ 1 $E_{\mathrm{pri}}=\Phi(S_{\mathrm{pri}})$;
2 $S_{0}\leftarrow$$\mathrm{RANDOM\_API}$ ($N_{\mathrm{syn}}$*$*\mathrm{L}$);
3 for iteration $t=0$ to $T-1$ do
@@ 6            $E_{t}=\Phi(S_{t})$
8             $S_{t}^{k}\leftarrow$$\mathrm{VARIATION\_API}$ $(S_{t})$ for $k=1,2\ldots,K$ $E_{t}=\frac{1}{K}\sum_{k=1}^{K}\Phi(S_{t}^{k})$
@@ 11       $P_{t}\leftarrow$ $\mathrm{Histogram}_{t}$ / sum ($\mathrm{Histogram}_{t}$);
14             $S_{t}^{\prime}\leftarrow$ draw $N_{\mathrm{syn}}$ samples with replacement from $S_{t}$ with probability $P_{t}$ $S_{t+1}\leftarrow$$\mathrm{VARIATION\_API}$$(S_{t}^{\prime})$ save dataset ${S_{\mathrm{syn}}}_{t+1}\leftarrow S_{t+1}$
@@
16             $S_{t}^{\prime}\leftarrow$ rank samples by probabilities $P_{t}$ and draw top $N_{\mathrm{syn}}$ samples save dataset ${S_{\mathrm{syn}}}_{t+1}\leftarrow S_{t}^{\prime}$ $S_{t+1}^{j}\leftarrow\mathrm{VARIATION\_API}(S_{t}^{\prime})$ for $\mathbf{j=1,2\ldots,L-1}$ $S_{t+1}\leftarrow[S_{t+1}^{1},...,S_{t+1}^{L-1},\mathbf{S_{t}^{\prime}}]$
17
18return ${S_{\mathrm{syn}}}_{T}$ Procedure DP_NN_HISTOGRAM($E_{\mathrm{syn}},E_{\mathrm{pri}},\sigma$)
@@
19       $\mathrm{Histogram}\leftarrow[0,...,0]$ for  $e_{\mathrm{pri}}\in E_{\mathrm{pri}}$  do
20             $i=\arg\min_{j\in[n]}d(e_{\mathrm{pri}},e_{j})$; $\mathrm{Histogram}[i]\leftarrow\mathrm{Histogram}[i]+1$
21      $\mathrm{Histogram}\leftarrow\mathrm{Histogram}+\mathcal{N}(0,\sigma^{2}I_{n})$ return $\mathrm{Histogram}$
@@ 4.1 Understanding the Performance of Aug-PE
Here, we analyze the performance of Aug-PE by answering four research questions about its utility and efficiency under DP compared to DP-finetuning-based baselines.
@@
RQ 3: How does Aug-PE perform across different privacy budget $\epsilon$? (1) Tb. 1 shows that Aug-PE in general achieves better performance as $\epsilon$ increases from $1,2,4$ to $\infty$, suggesting that Aug-PE scales well with the privacy budget $\epsilon$. (2) On OpenReview, from $\epsilon=\infty\rightarrow 1$, the rating classification accuracy obtained from DP-FT-Generator GPT-2-Large generated text drops from $0.4828\rightarrow 0.3855$, and DP-FT-Downstream on full training data drops from $0.6515\rightarrow 0.3052$, while the accuracy of Aug-PE GPT-3.5 exhibits marginal drop $0.4536\rightarrow 0.431$. It suggests that in some cases, the performance of Aug-PE (paired with powerful generator) can be more robust under DP noise than FT baselines. The reason could be that LMs are vulnerable to the perturbations introduced in model parameters through DP-SGD, whereas Aug-PE strategically adds noise to the histogram votes, effectively preserving the utility.
@@
GPT-3.5 67.9 74.7 66.8 74.7 45.4 43.5 41.9 43.1 30.4 32.7 30.1 32.4
RQ 4: Compared to DP-FT-Generator, how efficient the API-access-based Aug-PE is in terms of GPU hours? With inference API access, Aug-PE is more efficient than DP-FT-Generator that requires DP-SGD finetuning. (1) As shown in Fig. 6, to generate 100k synthetic samples on Yelp under $\epsilon=1$, given the same generator GPT-2-Large, Aug-PE $L=7$ provides 12.7x speedup and $L=2$ further provides 65.7x speedup. (2) The running time of Aug-PE is mainly scaled with # API calls, which is associated with the number of variations $L-1$ in Alg. 1. (3) The bottleneck of DP-FT-Generator is DP-SGD finetuning: it takes 1764 GPU hours on 32G NVIDIA V100 to finetune GPT-2-Large on Yelp and 7 hours to generate 100k samples, while Aug-PE $L=2$ ($L=7$) only requires 27 hours (139 hours). It highlights the computational expense of DP-SGD training, particularly for training LLMs, and underscores the efficiency of the API-based DP algorithm Aug-PE. A detailed breakdown of the GPU hours for each setting is in Appendix Tb. 22. (4) We use half precision (FP16) for LLM inference in Aug-PE. With the emerging efficient inference techniques (e.g., Liu et al. (2023)), Aug-PE runtime can be further optimized.
@@ 4.2 Understanding the Properties of Aug-PE
RQ 5: Can Aug-PE produce sentence length distributions similar to real data? Aug-PE produces favorable text length distributions. From Fig. 6, we see that the text length distribution of synthetic samples produced from GPT-3.5 through Aug-PE is close to the distribution of the original Yelp data, highlighting the effectiveness of our adaptive sequence length mechanism (§ 3.2). Note that the finetuning baseline requires a fixed $\mathrm{max\_token}$ (e.g., 128 for GPT-2), which leads to a hard threshold for maximal text length, which is not the case in our method with our adaptive length technique. Nevertheless, there is a peak near 30 tokens for Aug-PE, which is due to the $\mathrm{\min\_word}$ set in the prompt to prevent empty generation. We defer the convergence of text length distributions over PE iterations to § C.1.
RQ 6: Can Aug-PE benefit from more powerful LLMs? Aug-PE is effective across a wide range of API-accessible LLMs. We have observed from Tb. 1 that GPT-3.5 can lead to higher downstream accuracy than GPT-2-series, especially on PubMed and OpenReview. Here we evaluate more API-accessible, non-GPT based LLMs. (1) As shown in Tb. 2, under $\epsilon=\infty,1$, those modern LLMs can obtain comparable and even higher accuracy than GPT-3.5 on Yelp, suggesting that Aug-PE can effectively elicit and select high-quality synthetic text from various types of LLMs. Note that DP finetuning often needs to be implemented case-by-case for LLMs and currently lacks open-source implementations for these LLMs, whereas Aug-PE can easily leverage them. (2) The results on OpenReview and PubMed in Tb. 2 show that GPT-3.5 leads to higher utility than opensource LLMs (e.g. LLaMA-2), demonstrating the stronger generation power of GPT-3.5 in academic/medical domains. Interestingly, Mixtral-8x7B can also generate high-quality synthetic texts for OpenReview, but not for PubMed.
@@
RQ 7: Can more powerful downstream models benefit from synthetic text generated via Aug-PE? The high-quality synthetic text from Aug-PE is better utilized by larger downstream models. (1) From each row in Tb. 3, we see that next-word prediction accuracy monotonically increases with the use of larger downstream models trained on PubMed synthetic text. (2) Under both $\epsilon=1,\infty$, the smallest model $\text{BERT}_{\text{Tiny}}$ favors the synthetic texts from DP-FT-Generator GPT-2-Large, while larger models such as LLaMA-2 favor synthetic text from Aug-PE GPT-3.5. This observation underscores the importance of choosing downstream models of a suitable size; employing overly small models could under-estimate the quality of synthetic texts produced by Aug-PE with GPT-3.5. We hypothesize that this is because i) GPT-3.5 generated texts might already be of higher quality in terms of vocabulary, syntax, semantic coherence, etc., compared to generated texts from finetuned GPT-2-Large; and ii) larger downstream LMs like LLaMA-2 can better understand and utilize the nuances in synthetic texts for improved performance than $\text{BERT}_{\text{Tiny}}$.
RQ 8: Can we further improve downstream task accuracy with more synthetic samples generated from Aug-PE? To study the scaling law of Aug-PE, we use GPT-2-series models to generate {5k,10k,100k} samples for Yelp, and {2k,3k,5k} samples for other two datasets. As shown in § C.8, under $\epsilon=1,2,4,\infty$, Aug-PE in general achieves better performance across all datasets as the data size increases, suggesting that Aug-PE scales well with the number of synthetic samples.
@@ 4.3 Validating the Design of Aug-PE
RQ 9: Can Aug-PE surpass original PE? Tb. 4 shows that Aug-PE achieves notable improvement over PE for GPT-2, e.g., +22.6% on Yelp rating classification. We observe similar conclusions for GPT-3.5 in Tb. 23 in App.
@@
RQ 10: How does the private data guided sample selection affect Aug-PE performance? Here we aim to verify the components related to sample selection: i) usage of private data; ii) rank-based selection; iii) embedding model used during nearest neighbor voting.
i) Usage of private data. Tb. 5 shows that the initial samples (generated from Random API) or their variants (generated from Random API + Variation API) exhibit limited utility without using private data. However, the quality of the synthetic text improves notably after just one iteration of Aug-PE ($t=1$) when guided by private data, and this improvement continues to amplify with $T$ iterations.
@@
iii) Embedding models. Tb. 6 shows that larger embedding models such as “sentence-t5-xl” can more accurately capture the nuances of texts in the embedding space, leading to higher utility for GPT-2 generated texts.
@@
i) Variation API prompt designs. We evaluate the impact of four types of Variation API prompts on Yelp: paraphrasing and fill-in-the-blanks prompts under zero-shot and few-shot settings. (1) Qualitatively, we observed that GPT-2 struggles to adhere to the fill-in-the-blanks instruction, often leaving blanks (“____\_\__ _”) in the generated texts. In contrast, GPT-3.5 can effectively fill in the blanks, potentially because GPT-3.5 has been instruction-tuned (Wei et al., 2021) and thus follows the instructions better. (2) The quantitative results in Appendix Tb. 24 reveal that paraphrasing can be an effective strategy for GPT-2, while fill-in-the-blanks yields better results for GPT-3.5. (3) Fill-in-the-blanks offers more control over the diversity of generated content. By increasing the mask probability $p\%$, we can create more room for imaginative responses from GPT-3.5, leading to more diverse generations. As indicated in Fig. 6, a higher mask probability corresponds to increased accuracy in downstream area classification tasks when using GPT-3.5.
@@
ii) Temperature is a key parameter in controlling the diversity of LLM generation. A higher temperature leads LLMs to generate less frequent tokens, thereby increasing diversity. However, an excessively high temperature may result in overly random outputs and potentially hurt generation. The impact of different temperatures for Aug-PE on GPT-2 is shown in Tb. 7. (1) On Yelp, a higher temperature (1.4 to 1.7) proves beneficial for GPT-2, as business reviews often encompass daily conversations with a variety of sentence formats and tones. Additional findings in Fig. 6 indicate that large temperatures can also lead to low (better) FID scores for GPT-3.5. (2) Conversely, on OpenReview and PubMed, a moderate temperature setting (around 1.0) is more suitable for GPT-2, as academic and medical literature demand more precise and accurate text generation.
@@
iii) Increasing the number of variations $L-1$ generally enhances performance of Aug-PE as shown in Tb. 8, due to the expansion of the candidate synthetic sample pool, which increases the likelihood of getting high-quality texts. However, generating more variations requires additional API calls, leading to increased computational costs as discussed in Fig. 6. To balance the trade-off between utility and efficiency, we use $L=7$ for GPT-2-series experiments.

### v2

@@ 1 Introduction
Figure 1: Instead of finetuning LLMs with DP-SGD to generate synthetic text, Aug-PE only requires inference APIs of LLMs. Aug-PE works with the latest open-source LLMs and API-based LLMs to generate DP synthetic text with improved utility on OpenReview dataset, where DP-SGD finetuning is either hard to implement or infeasible.
The state-of-the-art DP synthetic text approach is to finetune pretrained generative language models (LMs) on private data with DP-SGD (a DP variant of SGD (Abadi et al., 2016)(Yue et al., 2023; Kurakin et al., 2023; Mattern et al., 2022a) (short-handed as DP finetune generator; see Fig. 1). Unlike non-DP ML applications, which have been greatly advanced by powerful LLMs such as GPT-4 (OpenAI, 2023b) and LLaMA (Touvron et al., 2023a, b) in a short time after they are released, the state-of-the-art DP synthetic text approaches are unfortunately still based on GPT-2.111The 175 billion-parameter GPT-3 has also been used for DP synthetic text (He et al., 2022). However, the solution is not publicly accessible as GPT-3 is proprietary. The reasons are: (1) Many powerful LLMs such as GPT-4, Claude, and Bard are only accessible through APIs. DP finetuning them is not feasible. [^2] Although standard finetuning APIs are provided for some of the models (OpenAI, 2023a), DP finetuning requires a special implementation and no model provides this custom API to date. (2) Even though some LLMs (e.g., LLaMA) are open-source, finetuning them with DP is resource-intensive and non-trivial to implement due to the need to calculate per-sample gradients (see § 2).
@@ 2 Background
DP synthetic text. To guarantee DP for private training data, one method involves using DP-SGD (Abadi et al., 2016) during model training for specific NLP tasks (Yu et al., 2022; Li et al., 2021). Alternatively, one can finetune pretrained generative language models, such as GPT-2, with private data using DP-SGD and then generate synthetic text datasets (Putta et al., 2022; Bommasani et al., 2019) (Fig. 1). Such DP synthetic texts can be employed in an arbitrary number of non-privately trained downstream tasks without increasing privacy loss. Studies by Yue et al. (2023); Mattern et al. (2022a); Kurakin et al. (2023) indicate that training downstream models on DP synthetic text yields performance akin to directly training them on real data with DP, highlighting the good quality of synthetic data.
@@
Additionally, there is a line of work on text-to-text privatization techniques, which provide different privacy guarantees than DP, such as word-level metric DP or sample-level local DP. We defer more discussion and comparison to § C.11.
@@ 1 $E_{\mathrm{pri}}=\Phi(S_{\mathrm{pri}})$;
2
[…]
4
5for iteration $t=0$ to $T-1$ do
6
@@ 9            $E_{t}=\Phi(S_{t})$
11             $S_{t}^{k}\leftarrow$$\mathrm{VARIATION\_API}$ $(S_{t})$ for $k=1,2\ldots,K$
12            $E_{t}=\frac{1}{K}\sum_{k=1}^{K}\Phi(S_{t}^{k})$
@@ 16      $P_{t}\leftarrow$ $\mathrm{Histogram}_{t}$ / sum ($\mathrm{Histogram}_{t}$);
20             $S_{t}^{\prime}\leftarrow$ draw $N_{\mathrm{syn}}$ samples with replacement from $S_{t}$ with probability $P_{t}$
21            $S_{t+1}\leftarrow$$\mathrm{VARIATION\_API}$$(S_{t}^{\prime})$
22            save dataset ${S_{\mathrm{syn}}}_{t+1}\leftarrow S_{t+1}$
@@ 21            $S_{t+1}\leftarrow$$\mathrm{VARIATION\_API}$$(S_{t}^{\prime})$
24             $S_{t}^{\prime}\leftarrow$ rank samples by probabilities $P_{t}$ and draw top $N_{\mathrm{syn}}$ samples
25            save dataset ${S_{\mathrm{syn}}}_{t+1}\leftarrow S_{t}^{\prime}$
26            $S_{t+1}^{j}\leftarrow\mathrm{VARIATION\_API}(S_{t}^{\prime})$ for $\mathbf{j=1,2\ldots,L-1}$
27            $S_{t+1}\leftarrow[S_{t+1}^{1},...,S_{t+1}^{L-1},\mathbf{S_{t}^{\prime}}]$
28
29return ${S_{\mathrm{syn}}}_{T}$
30 Procedure DP_NN_HISTOGRAM($E_{\mathrm{syn}},E_{\mathrm{pri}},\sigma$)
@@
31
32      $\mathrm{Histogram}\leftarrow[0,...,0]$
33      for  $e_{\mathrm{pri}}\in E_{\mathrm{pri}}$  do
34             $i=\arg\min_{j\in[n]}d(e_{\mathrm{pri}},e_{j})$;
35            $\mathrm{Histogram}[i]\leftarrow\mathrm{Histogram}[i]+1$
36      $\mathrm{Histogram}\leftarrow\mathrm{Histogram}+\mathcal{N}(0,\sigma^{2}I_{n})$
37      return $\mathrm{Histogram}$
@@ 4.1 Understanding the Performance of Aug-PE
Here, we analyze the performance of Aug-PE by answering five research questions about its utility, efficiency, and robustness against empirical privacy attacks under DP compared to DP-finetuning-based baselines.
@@
RQ 3: How does Aug-PE perform across different privacy budget $\epsilon$? (1) Tb. 1 shows that Aug-PE in general achieves better performance as $\epsilon$ increases from $1,2,4$ to $\infty$, suggesting that Aug-PE scales well with the privacy budget $\epsilon$. (2) On OpenReview, from $\epsilon=\infty\rightarrow 1$, the rating classification accuracy obtained from DP-FT-Generator GPT-2-Large generated text drops from $48.3\rightarrow 38.6$, and DP-FT-Downstream on full training data drops from $65.1\rightarrow 30.5$, while the accuracy of Aug-PE GPT-3.5 exhibits marginal drop $45.4\rightarrow 43.1$. It suggests that in some cases, the performance of Aug-PE (paired with powerful generator) can be more robust under DP noise than FT baselines. The reason could be that LMs are vulnerable to the perturbations introduced in model parameters through DP-SGD, whereas Aug-PE strategically adds noise to the histogram votes, effectively preserving the utility.
@@
GPT-3.5 68.4 74.1 67.9 74.0 45.4 43.5 41.9 43.1 30.4 32.7 30.1 32.4
RQ 4: Compared to DP-FT-Generator, how efficient the API-access-based Aug-PE is in terms of GPU hours? With inference API access, Aug-PE is more efficient than DP-FT-Generator that requires DP-SGD finetuning. (1) As shown in Fig. 6, to generate 100k synthetic samples on Yelp under $\epsilon=1$, given the same generator GPT-2-Large, Aug-PE $L=7$ provides 12.7x speedup and $L=2$ further provides 65.7x speedup. (2) The running time of Aug-PE is mainly scaled with # API calls, which is associated with the number of variations $L-1$ in Alg. 1. (3) The bottleneck of DP-FT-Generator is DP-SGD finetuning: it takes 1764 GPU hours on 32G NVIDIA V100 to finetune GPT-2-Large on Yelp and 7 hours to generate 100k samples, while Aug-PE $L=2$ ($L=7$) only requires 27 hours (139 hours). It highlights the computational expense of DP-SGD training, particularly for training LLMs, and underscores the efficiency of the API-based DP algorithm Aug-PE. A detailed breakdown of the GPU hours for each setting is in Appendix Tb. 23. (4) We use half precision (FP16) for LLM inference in Aug-PE. With the emerging efficient inference techniques (e.g., Liu et al. (2023)), Aug-PE runtime can be further optimized.
RQ 5: How robust Aug-PE is under empircal privacy attacks compared to DP-finetuning-based baselines? We perform state-of-the-art text membership inference attacks (MIAs) against the finetuned downstream models on PubMed dataset. We consider three types of MIAs and report the AUC score: (1) PPL thresholds perplexity to predict membership (Carlini et al., 2021); (2) REFER computes the ratio of the log perplexity of the tested model against a reference model (Carlini et al., 2021); (3) LIRA uses the ratio of likelihood (Carlini et al., 2022) and we use the pre-trained model as a reference following (Mattern et al., 2023). The results in Tb. 3 show that Aug-PE generally exhibits lower AUC scores under MIAs compared to DP-FT-Generator and DP-FT-Downstream. This indicates a higher robustness to empirical privacy attacks, potentially due to the synthetic nature of the data used for downstream model finetuning, which inherently reduces the risk of overfitting to real private data. We defer the details to § C.1.
Table 3: Aug-PE generally yields lower AUC scores against membership inference attacks on PubMed than DP-FT-Generator and DP-FT-Downstream, indicating a higher robustness to empirical privacy attacks.
Method Generator AUC ($\epsilon=\infty$) AUC ($\epsilon=4$) AUC ($\epsilon=2$) Avg
PPL REFER LIRA PPL REFER LIRA PPL REFER LIRA
DP-FT-Downstream / 77.60 74.93 65.05 49.32 54.58 56.82 48.96 50.41 51.56 58.80
DP-FT-Generator GPT-2 55.50 51.35 51.97 53.90 51.12 51.84 53.31 50.81 51.61 52.38
DP-FT-Generator GPT2-M 54.91 51.25 51.88 54.72 51.13 51.76 54.58 51.28 51.85 52.60
DP-FT-Generator GPT2-L 54.81 51.22 51.86 54.56 50.81 51.64 55.05 51.01 51.69 52.52
Aug-PE GPT-2 50.08 50.92 51.66 50.10 50.97 51.70 49.94 50.85 51.64 50.87
Aug-PE GPT2-M 49.85 50.73 51.57 50.10 50.95 51.69 49.73 50.65 51.51 50.75
Aug-PE GPT2-L 49.43 50.40 51.40 49.61 50.56 51.48 49.66 50.60 51.49 50.51
Aug-PE GPT-3.5 52.23 49.67 50.85 52.68 49.84 50.93 52.77 49.77 50.85 51.07
@@ 4.2 Understanding the Properties of Aug-PE
RQ 6: Can Aug-PE produce sentence length distributions similar to real data? Aug-PE produces favorable text length distributions. From Fig. 6, we see that the text length distribution of synthetic samples produced from GPT-3.5 through Aug-PE is close to the distribution of the original Yelp data, highlighting the effectiveness of our adaptive sequence length mechanism (§ 3.2). Note that the finetuning baseline requires a fixed $\mathrm{max\_token}$ (e.g., 128 for GPT-2), which leads to a hard threshold for maximal text length, which is not the case in our method with our adaptive length technique. Nevertheless, there is a peak near 30 tokens for Aug-PE, which is due to the $\mathrm{\min\_word}$ set in the prompt to prevent empty generation. We defer the convergence of text length distributions over PE iterations to § C.2.
RQ 7: Can Aug-PE benefit from more powerful LLMs? Aug-PE is effective across a wide range of API-accessible LLMs. We have observed from Tb. 1 that GPT-3.5 can lead to higher downstream accuracy than GPT-2-series, especially on PubMed and OpenReview. Here we evaluate more API-accessible, non-GPT based LLMs. (1) As shown in Tb. 2, under $\epsilon=\infty,1$, those modern LLMs can obtain comparable and even higher accuracy than GPT-3.5 on Yelp, suggesting that Aug-PE can effectively elicit and select high-quality synthetic text from various types of LLMs. Note that DP finetuning often needs to be implemented case-by-case for LLMs and currently lacks open-source implementations for these LLMs, whereas Aug-PE can easily leverage them. (2) The results on OpenReview and PubMed in Tb. 2 show that GPT-3.5 leads to higher utility than opensource LLMs (e.g. LLaMA-2), demonstrating the stronger generation power of GPT-3.5 in academic/medical domains. Interestingly, Mixtral-8x7B can also generate high-quality synthetic texts for OpenReview, but not for PubMed.
@@
RQ 8: Can more powerful downstream models benefit from synthetic text generated via Aug-PE? The high-quality synthetic text from Aug-PE is better utilized by larger downstream models. (1) From each row in Tb. 4, we see that next-word prediction accuracy monotonically increases with the use of larger downstream models trained on PubMed synthetic text. (2) Under both $\epsilon=1,\infty$, the smallest model $\text{BERT}_{\text{Tiny}}$ favors the synthetic texts from DP-FT-Generator GPT-2-Large, while larger models such as LLaMA-2 favor synthetic text from Aug-PE GPT-3.5. This observation underscores the importance of choosing downstream models of a suitable size; employing overly small models could under-estimate the quality of synthetic texts produced by Aug-PE with GPT-3.5. We hypothesize that this is because i) GPT-3.5 generated texts might already be of higher quality in terms of vocabulary, syntax, semantic coherence, etc., compared to generated texts from finetuned GPT-2-Large; and ii) larger downstream LMs like LLaMA-2 can better understand and utilize the nuances in synthetic texts for improved performance than $\text{BERT}_{\text{Tiny}}$.
RQ 9: Can we further improve downstream task accuracy with more synthetic samples generated from Aug-PE? To study the scaling law of Aug-PE, we use GPT-2-series models to generate {5k,10k,100k} samples for Yelp, and {2k,3k,5k} samples for other two datasets. As shown in § C.10, under $\epsilon=1,2,4,\infty$, Aug-PE in general achieves better performance across all datasets as the data size increases, suggesting that Aug-PE scales well with the number of synthetic samples.
@@ 4.3 Validating the Design of Aug-PE
RQ 10: Can Aug-PE surpass original PE? Tb. 5 shows that Aug-PE achieves notable improvement over PE for GPT-2, e.g., +22.6% on Yelp rating classification. We observe similar conclusions for GPT-3.5 in Tb. 24 in App.
@@
RQ 11: How does the private data guided sample selection affect Aug-PE performance? Here we aim to verify the components related to sample selection: i) usage of private data; ii) rank-based selection; iii) embedding model used during nearest neighbor voting.
i) Usage of private data. Tb. 6 shows that the initial samples (generated from Random API) or their variants (generated from Random API + Variation API) exhibit limited utility without using private data. However, the quality of the synthetic text improves notably after just one iteration of Aug-PE ($t=1$) when guided by private data, and this improvement continues to amplify with $T$ iterations. We report the results under DP in § C.8.
@@
iii) Embedding models. Tb. 7 shows that larger embedding models such as “sentence-t5-xl” can more accurately capture the nuances of texts in the embedding space, leading to higher utility for GPT-2 generated texts.
@@
i) Variation API prompt designs. We evaluate the impact of four types of Variation API prompts on Yelp: paraphrasing and fill-in-the-blanks prompts under zero-shot and few-shot settings. (1) Qualitatively, we observed that GPT-2 struggles to adhere to the fill-in-the-blanks instruction, often leaving blanks (“____\_\__ _”) in the generated texts. In contrast, GPT-3.5 can effectively fill in the blanks, potentially because GPT-3.5 has been instruction-tuned (Wei et al., 2021) and thus follows the instructions better. (2) The quantitative results in Appendix Tb. 25 reveal that paraphrasing can be an effective strategy for GPT-2, while fill-in-the-blanks yields better results for GPT-3.5. (3) Fill-in-the-blanks offers more control over the diversity of generated content. By increasing the mask probability $p\%$, we can create more room for imaginative responses from GPT-3.5, leading to more diverse generations. As indicated in Fig. 6, a higher mask probability corresponds to increased accuracy in downstream area classification tasks when using GPT-3.5.
@@
ii) Temperature is a key parameter in controlling the diversity of LLM generation. A higher temperature leads LLMs to generate less frequent tokens, thereby increasing diversity. However, an excessively high temperature may result in overly random outputs and potentially hurt generation. The impact of different temperatures for Aug-PE on GPT-2 is shown in Tb. 8. (1) On Yelp, a higher temperature (1.4 to 1.7) proves beneficial for GPT-2, as business reviews often encompass daily conversations with a variety of sentence formats and tones. Additional findings in Fig. 6 indicate that large temperatures can also lead to low (better) FID scores for GPT-3.5. (2) Conversely, on OpenReview and PubMed, a moderate temperature setting (around 1.0) is more suitable for GPT-2, as academic and medical literature demand more precise and accurate text generation.
@@
iii) Increasing the number of variations $L-1$ generally enhances performance of Aug-PE as shown in Tb. 9, due to the expansion of the candidate synthetic sample pool, which increases the likelihood of getting high-quality texts. However, generating more variations requires additional API calls, leading to increased computational costs as discussed in Fig. 6. To balance the trade-off between utility and efficiency, we use $L=7$ for GPT-2-series experiments.